OpenStudy (anonymous):

Carbon dioxide in a soft drink with an initial mass of \(381.04\)g was carefully extracted and collected as a gas. The drink final mass was \(380.41\)g. What would be the volume that the Carbon Diodixide occupy at \(100\)kPa and 25 degrees celsius?

OpenStudy (anonymous):

Anyone able to give me a hint? what formula do i use? I haven't done these in a long time..

OpenStudy (anonymous):

ok so you have mass of the co2 dont you? you can do this two ways, one assumes that conditions (temperature and pressure) are ideal or near ideal so you calculate from: n=m/M n=V/Vm -> V = n*Vm or the other way is to calculate the conditions in and use: pV=nRT -> V = nRT/p where again n=m/M

OpenStudy (anonymous):

Im not getting it \[n=\frac{m}{M} => n = \frac{381.04}{268} = 1.4217\] \[n=\frac{V}{V_{m}} => 1.4217 = \frac{380.41}{V} => V = 267..\] where did i go wrong?

OpenStudy (anonymous):

381,04-380,41 is the mass!

OpenStudy (anonymous):